# Deformation dynamics of an oil droplet into a crescent shape during the intermittent motion

### v1

@@ IV Discussion
IV.1 Quantitative model for droplet deformation
The motion and deformation of a droplet are attributed to the surface tension difference induced by camphor molecules. However, a large deformation of a droplet shape without fission has not been observed without oil red O[40]. This suggests that the presence of oil red O molecules decreases the interfacial tension between water and oil, and this enables a large deformation. As illustrated in Figs. 6 and 7, the equilibration of the oil/water interface in the presence of oil red O molecules is slow since the characteristic time for the decrease in interfacial tension is around 800 s. Moreover, the camphor concentration in the droplet decreases over time. As a result, the gradient of interfacial tension around the droplet gets smaller in time, and the character of motion changes from translational to intermittent with dynamic deformation into a crescent shape.
Now, let us present quantitative arguments on how the decrease in interfacial tension induced deformation by evaluating the energy of deformation. If we neglect the contribution of the periphery, then the free energy of a floating droplet shown in Fig. 8 is estimated as:
@@ IV.1 Quantitative model for droplet deformation
Here, $S=\gamma_{\mathrm{aw}}-(\gamma_{\mathrm{ow}}+\gamma_{\mathrm{ao}})$ is the spreading coefficient, $\gamma_{\mathrm{aw}},\gamma_{\mathrm{ow}}$ and $\gamma_{\mathrm{ao}}$ are the interfacial tensions between air and water, oil and water, and air and oil, respectively, $\Sigma$ is the area of the droplet, and $\Delta\rho_{\mathrm{ao}}$ and $\Delta\rho_{\mathrm{ow}}$ are the density differences between air and oil and that between oil and water, respectively, as shown in Fig. 8. $e_{\mathrm{ao}}$ and $e_{\mathrm{ow}}$ are the differences in height between the water level at infinity and the air/oil interface, and that between the water level at infinity and the oil/water interface, respectively. We neglected the second term as it is much smaller than the first term since the thickness of the droplet is small.
@@
Additionally, a droplet should have excess energy since its periphery is curved. The excess energy per unit periphery length equals
@@
The detail in the derivation is in Appendix B. The free energy of droplet deformation is described as:
$\Delta E=\Delta E_{\Sigma}+\Delta E_{l},$ (6)
where $\Delta E_{\Sigma}=-S\Delta\Sigma$ is the energy for an increase in droplet area $\Delta\Sigma$, and $\Delta E_{l}=\mathcal{T}\Delta l$ is energy for an increase in droplet periphery $\Delta l$. If the surface tension difference between the front and back of the droplet is $\delta\gamma_{\mathrm{aw}}$ and the droplet swept a water surface with an area of $\sigma$, then the work exerted on the droplet is estimated as
@@
Figure 9: Dependence of energy of deformation on the interfacial tension between oil and water. (a) Plots of $\Delta E_{\Sigma}$ (red), $\Delta E_{l}$ (orange) and $W$ (blue), respectively. (b) Energy of deformation normalized by the work exerted on the droplet, $\Delta E/W$ against $\gamma_{\mathrm{ow}}$.
IV.2 Deformation of a droplet with oil red O
The time evolution of the interfacial tension shown in Fig. 6 and the discussion in the energy of deformation in the previous subsection indicate that the droplet deformation becomes easier with time $t$ after it was placed on the water surface. To confirm this, we investigated the deformation of a camphor-free droplet with oil red O by touching a camphor disk to the water surface close to the droplet. We put a 20 $\mu$L droplet of paraffin oil with oil red O (concentration 1.0 g/L) and without camphor on the water surface. Then, after time $T$, we touched a 3 mm-diameter camphor disk to the water surface at around 5 mm from the periphery of the droplet. Then, the droplet escaped from the camphor disk with deformation into a crescent shape as shown in Fig. 10(a). We performed experiments for $T=180,300,600$ s and plotted the characteristic parameters describing the droplet shape dynamics in Fig. 10(b), which are defined in the same manner as in Sec. III. The time $\Delta T=0$ corresponds to when the droplet COM speed $v$ has the maximum value. The maximum magnitude of deformation $A_{\mathrm{max}}$ increased with $T$ as shown in Fig. 10(b-2). The shape of the droplet was greatly curved when $T=600$ s, which should be the reason why $R_{\mathrm{max}}$ was smaller than that when $T=300$ s. Figure 10(b-3) shows a scatter plot of the times $\Delta T_{R}$ and $\Delta T_{A}$, at which the maximum values of $R$ and $A$ ($R_{\mathrm{max}}$ and $A_{\mathrm{max}}$) were observed. Both $\Delta T_{R}$ and $\Delta T_{A}$ were positive and had almost the same value. These behaviors were the same as those for the spontaneous deformation of a droplet with camphor shown in Fig. 4, though the relationship between $\Delta T_{R}$ and $\Delta T_{A}$ was slightly different from that between $\Delta t_{R}$ and $\Delta t_{A}$. This may be due to the much larger deformation in the experiments with a camphor-free droplet. When $T=30$ s, the droplet hardly deformed as shown in Fig. 10(c). The result in Fig. 10 shows that for longer $T$ droplet deformation becomes easier.
Figure 10: Deformation of a droplet (oil red O concentration 1.0 g/L) observed when the water surface was touched with a camphor disk after elapsed time $T$ since the droplet had been placed on the water surface. The time $\Delta T=0$ corresponds to when $v$ has the maximum value. (a) Snapshots of a droplet for $T=300$ s. (b) The characteristic values for the droplet deformation. (b-1) and (b-2) show the average value for $R_{\mathrm{max}}$ and $A_{\mathrm{max}}$ with the standard deviation as error bars for 3 experiments for each $T$, and (b-3) is a scatter plot of $\Delta T_{A}$ and $\Delta T_{R}$. (c) Superposed images of a droplet for $T=30$ s with a time difference of 2.0 s. The red line indicates the trajectories of the droplet COM.
@@ V Conclusions
A paraffin droplet containing camphor and oil red O dye exhibits spontaneous motion with deformation. At considered concentrations of camphor and oil red O, the droplet started to perform intermittent motion approximately 2000 s after placing it on the water surface. In the presented study, we introduced a quantitative model that describes the possibility of large deformation of a droplet during intermittent motion. The model predicts that the shape deformation occurs when the oil/water interfacial tension becomes lower. It can spontaneously happen for a droplet located on the water surface because, after the formation of the interface, oil/water interfacial tension gradually decreases in time due to the adsorption of oil red O at the oil/water interface. Simultaneously, the camphor concentration in the droplet also decreases with time, resulting in a smaller driving force. The combination of decreased interfacial tension due to oil red O and reduced driving force due to reduced camphor concentration in the droplet causes intermittent motion with dynamic deformation into a crescent shape.
@@ A.2 Analysis method
Appendix B Derivation of the equation for line tension
Assuming the characteristic length of deformation of a triple line is much larger than the capillary length, we consider the deformation of the interface near the straight triple line. We set the Cartesian coordinates so that the $y$-axis matches the triple line and $z=\zeta_{\mathrm{ao}}(x)(x>0)$ and $z=\zeta_{\mathrm{ow}}(x)(x>0)$ indicate the interface between air and oil and that between oil and water, respectively, as shown in Fig. 9. Then, the excess energy of the triple line per unit length is described as
@@ Appendix B Derivation of the equation for line tension
$+\qquad\frac{1}{2}\rho_{\mathrm{ow}}g(e_{\mathrm{ow}}-\zeta_{\mathrm{ow}}(x))^{2}\biggr{\}},$ (18)
@@
where $\kappa_{\mathrm{ao}}^{-1}$ and $\kappa_{\mathrm{ow}}^{-1}$ are capillary lengths of air/oil and oil/water interfaces, respectively. From these equations, we can derive Eq. (3).
Appendix C Water surface tension measurement around a droplet
We measured the surface tension around a 30 $\mu$L paraffin droplet with camphor (concentration 0.06 M) and oil red O (concentration 1.0 g/L) using quasi-elastic laser scattering (QELS) method[5]. We put a metal screw on the bottom of the Petri dish upside down and then placed a droplet on the tip of the screw so that the droplet was pinned. We measured the surface tension close to the spot where the droplet was pinned. The optical settings for the surface tension measurement were the same as described in [5].
The decrease in surface tension $\gamma_{\mathrm{aw}}-\gamma_{\mathrm{aw}}^{(0)}$ is plotted as a function of time $t$ in Fig. 12, in which the initial value $\gamma_{\mathrm{aw}}^{(0)}$ is determined by fitting the time series for $0~{}\mathrm{s}\leq t\leq 40~{}\mathrm{s}$ to an exponential function
$\gamma_{\mathrm{aw}}(t)=\gamma_{\mathrm{aw}}^{(0)}-\Delta\gamma_{\mathrm{aw}}(1-e^{-t/\tau}).$ (23)
Time $t=0$ corresponds to the time when the droplet was placed on the tip of the screw. The surface tension slightly decreased by $\Delta\gamma_{\mathrm{aw}}=0.24$ mN/m.
Figure 12: Decrease in the surface tension depending on time $t$ after putting a droplet on the water surface. The red line shows the exponential function obtained by the fitting with Eq. (23). The paraffin droplet included camphor at 0.06 M and oil red O at 1.0 g/L.

### v2

@@ IV Discussion
IV.1 Analysis of free energy related to droplet deformation
To better understand the droplet deformation phenomenon, we evaluated two related dimensionless numbers. First, the capillary number $\mathrm{Ca}$ is the ratio of the viscous force and interfacial tension, $\mathrm{Ca}=\mu v/\gamma_{\mathrm{ow}}$, where $\mu$ is the viscosity of the liquid, $v$ is the characteristic velocity, and $\gamma_{\mathrm{ow}}$ is the characteristic value for oil/water interfacial tension. By substituting $\mu\simeq 10^{-3}~{}\mathrm{Pa\,s}$, $v\simeq 10^{-2}~{}\mathrm{m/s}$, and $\gamma_{\mathrm{ow}}\simeq 60\times 10^{-3}~{}\mathrm{N/m}$, we obtain $\mathrm{Ca}\simeq 2\times 10^{-4}\ll 1$. This means that during an intermittent motion, the relative effect of viscous drag forces is much smaller than that of surface tension.
Next, the Weber number $\mathrm{We}$ is the relative importance of the fluid inertia compared to its surface tension. $\mathrm{We}=\rho rv^{2}/\gamma_{\mathrm{ow}}$, where $\rho$ is the mass density of the liquid and $r$ is the characteristic length. For $\rho\simeq 10^{3}~{}\mathrm{kg/m^{3}}$, $r\simeq 10^{-2}~{}\mathrm{m}$ and other parameter values listed above, we obtained $\mathrm{We}\simeq 2\times 10^{-4}\ll 1$. This means that the inertial effects are negligible during an intermittent motion. Considering both the dimensionless numbers, we concluded that the surface tension difference around the droplet plays the most important role in droplet deformation.
In the following, we use the approximation of droplet free energy to derive the energy condition for a spontaneous deformation of droplet shape. If we neglect the contribution of the periphery, then the free energy of a floating droplet shown in Fig. 8 is estimated as:
@@ IV.1 Analysis of free energy related to droplet deformation
Here, $S=\gamma_{\mathrm{aw}}-(\gamma_{\mathrm{ow}}+\gamma_{\mathrm{ao}})$ is the spreading coefficient, $\gamma_{\mathrm{aw}}$ and $\gamma_{\mathrm{ao}}$ are the interfacial tensions between air and water and between air and oil, respectively, $\Sigma$ is the area of the droplet, and $\Delta\rho_{\mathrm{ao}}$ and $\Delta\rho_{\mathrm{ow}}$ are the density differences between air and oil and that between oil and water, respectively, as shown in Fig. 8. In Eq. (2), $e_{\mathrm{ao}}$ and $e_{\mathrm{ow}}$ are the differences in height between the water level at infinity and the air/oil interface, and that between the water level at infinity and the oil/water interface, respectively. For the following analyses, we neglected the second term in Eq. (2) as it is much smaller than the first term since the thickness of a droplet is small.
@@
The second contribution to droplet energy comes from the curvature of its outer edge. The excess energy of the triple line per unit length equals:
@@
The detail in the derivation is presented in the Appendix A.3.
The change in free energy related to droplet deformation can be approximated as:
$\Delta E=-S\Delta\Sigma+\mathcal{T}\Delta l,$ (6)
where the first term is the energy for an increase in droplet area by $\Delta\Sigma$, and the second describes the energy for an increase in droplet periphery $\Delta l$. If the surface tension difference between the front and back of the droplet is $\delta\gamma_{\mathrm{aw}}$ and the droplet moved on a water surface by an area of $\sigma$, then the work exerted on the droplet is estimated as
@@
To support self-motion with deformation, the energy gain due to the surface tension gradient should exceed the free energy increase due to the droplet deformation depending on the interfacial tension. In our system, the quasi-elastic laser scattering (QELS) measurements shown in Appendix B suggested that the value of $\delta\gamma_{\mathrm{aw}}$ does not exceed $0.3~{}\mathrm{mN/m}$, though we were not able to measure this value directly. For the following calculations, we used $\delta\gamma_{\mathrm{aw}}=0.25~{}\mathrm{mN/m}$. During the intermittent motion shown in Fig. 3, we observed $\Delta\Sigma=3.3~{}\mathrm{mm}^{2}$, $\Delta l=17.2~{}\mathrm{mm}$ and $\sigma=1288.5~{}\mathrm{mm}^{2}$. For these parameters and values of interfacial tensions measured by the Wilhelmy method: $\gamma_{\mathrm{ao}}=31.6~{}\mathrm{mN/m}$ and $\gamma_{\mathrm{aw}}=73.5~{}\mathrm{mN/m}$, the values of $\Delta E_{\Sigma}=S\Delta\Sigma$, $\Delta E_{l}=\mathcal{T}\Delta l$ and $W$ are plotted as functions of $\gamma_{\mathrm{ow}}$ in Fig. 9(a). For these realistic values of parameters ($\gamma_{\mathrm{ow}}\sim 50~{}\mathrm{mN/m}$), $\Delta E_{\Sigma}+\Delta E_{l}$, and $W$ are of the same order, which implies that the deformation of a droplet can occur. $\Delta E_{\Sigma}(\gamma_{\mathrm{ow}})$ and $\Delta E_{l}(\gamma_{\mathrm{ow}})$ are increasing functions of interfacial tension $\gamma_{\mathrm{ow}}$, while $W$ does not depend on it. The increase in $\Delta E_{l}(\gamma_{\mathrm{ow}})$ is faster than that in $\Delta E_{\Sigma}(\gamma_{\mathrm{ow}})$, which means that the line energy is the most important for droplet deformation. Figure 9(b) shows the deformation energy normalized by the work exerted on the droplet. For the selected parameter values, the spontaneous droplet deformation can occur when $\Delta E/W<1$, which happens if $\gamma_{\mathrm{ow}}\lesssim 47~{}\mathrm{mN/m}.$ Having in mind that $\gamma_{\mathrm{ow}}$ is a decreasing function of oil red O concentration (cf. Fig. 6), we can see why droplet deformation occurs for droplets with a higher concentration of oil red O. The estimated threshold for deformation is a bit lower than expected from the results discussed in the previous chapter. There could be many reasons for this. First, the estimation is based on a simple model. Second, the local inhomogeneities in oil red O concentration, especially those near the droplet boundary, can trigger local shape changes. And finally, there may be possible errors in the measurements of interfacial tension. For example, the threshold should increase for the lower value of $\gamma_{\mathrm{ao}}$.
Figure 9: Major forms of energy related to droplet deformation as functions of the interfacial tension between oil and water. (a) Plots of $\Delta E_{\Sigma}$ (red), $\Delta E_{l}$ (orange) and $W$ (blue), respectively. (b) Deformation energy normalized by the work exerted on the droplet, $\Delta E/W$ against $\gamma_{\mathrm{ow}}$.
IV.2 Dynamics of droplet deformation induced by camphor release
The discussion of energetic aspects of droplet deformation indicates that a deformation occurs easier for a long time $t$ after the droplet is placed on the water surface. To verify this observation, we investigated deformations of paraffin droplets with oil red O that do not contain camphor. Of course, if such droplets are placed on the water surface they do not move and do not change their shapes. The surface tension gradient necessary for the motion was induced by touching the water surface close to the droplet with a camphor piece at time $T$ after the droplet had been placed on water. The experiment with a $20~{}\mu\mathrm{L}$ paraffin droplet containing oil red O (concentration $1.0~{}\mathrm{g/L}$) is illustrated in Fig. 10. As a camphor source, we used a $3~{}\mathrm{mm}$-diameter camphor disk located around $5~{}\mathrm{mm}$ away from the droplet boundary. After the camphor disk touched the surface, the droplet moved away from the source with deformation into a crescent shape, as shown in Fig. 10(a). We performed experiments for $T=30,180,300,600~{}\mathrm{s}$ and plotted the characteristic parameters describing the droplet shape dynamics in Fig. 10(b). The time $\Delta T=0$ corresponds to when the droplet COM speed $v$ has the maximum value. The maximum magnitude of deformation $A_{\mathrm{max}}$ increased with $T$ as shown in Fig. 10(b-2). The shape of the droplet was the most curved when $T=600~{}\mathrm{s}$, which can explain why $R_{\mathrm{max}}(T=600~{}\mathrm{s})$ was smaller than that when $T=300~{}\mathrm{s}$. Figure 10(b-3) shows a scatter plot of the times $\Delta T_{R}$ and $\Delta T_{A}$, at which the maximum values of $R$ and $A$ ($R_{\mathrm{max}}$ and $A_{\mathrm{max}}$) were observed. Both $\Delta T_{R}$ and $\Delta T_{A}$ were positive and had almost the same value. These behaviors were the same as those for the spontaneous deformation of a droplet with camphor shown in Fig. 4, though the relationship between $\Delta T_{R}$ and $\Delta T_{A}$ was slightly different from that between $\Delta t_{R}$ and $\Delta t_{A}$. This may be due to the much larger deformation in the experiments with a camphor-free droplet. When $T=30$ s, the droplet hardly deformed, as shown in Fig. 10(c), and the characteristic parameters could not be appropriately obtained. These results indicate that for longer $T$ of the order of $10$ minutes, the modification of the droplet interface becomes more efficient, leading to more pronounced droplet deformation.
Figure 10: Deformation of a paraffin droplet containing oil red O (concentration $1.0~{}\mathrm{g/L}$) observed when the water surface was touched with a camphor disk after elapsed time $T$ after the droplet had been placed on the water surface. The time $\Delta T=0$ corresponds to when $v$ has the maximum value. (a) Snapshots of a droplet for $T=300~{}\mathrm{s}$. (b) The values of parameters characterizing droplet deformation. Panels (b-1) and (b-2) show the average value for $R_{\mathrm{max}}$ and $A_{\mathrm{max}}$ with the standard deviation as error bars for 3 experiments for each $T$, and (b-3) is a scatter plot of $\Delta T_{A}$ and $\Delta T_{R}$. (c) Superposed images of a droplet for $T=30~{}\mathrm{s}$ with a time difference of $2.0~{}\mathrm{s}$. The red line indicates the trajectories of the droplet COM.
IV.3 Dynamics of flows around the droplet
In this subsection, we focus our attention on time-dependent velocity and deformation during the intermittent droplet motion. In particular, we focus our attention on the role of the oil red O because in the previous study, a large deformation of a droplet shape without fission has not been observed without oil red O[41].
The results shown in Figs. 3, 4 and 10 indicate complex interactions between camphor and oil red O. For a droplet forced to move by an external source, $\Delta T_{A}$ and $\Delta T_{R}$ seem perfectly correlated. In this case, the time difference between the peak of velocity and that of the deformation can be as long as $4~{}\mathrm{s}$ and increases with time $T$. For a droplet containing camphor, $\Delta t_{A}$ and $\Delta t_{R}$ are much shorter ($<1.5~{}\mathrm{s}$). Moreover, $\Delta t_{R}<\Delta t_{A}$, but the values are weakly correlated. The different relationship between the peaks can be attributed to the presence of surface active molecules inside a droplet.
@@ IV.3 Dynamics of flows around the droplet
The snapshots illustrating the evolution of droplets and flows are shown in Fig. 11. Before the droplet started the intermittent motion, the particles were still on the water surface, which means that no strong surface flow was induced around the droplet (see Fig. 11(a)). The evolution towards intermittent motion should be ruled by fluctuations in camphor release. A small region without tracer particles can be seen up-left of the droplet. The tracer particle-free area grows in time as seen in Figs. 11(b)–(c) and generates a difference in surface tension between the upper and the lower boundaries of the droplet. The resulting force is large enough to drive the droplet towards the region with low camphor concentration. After more camphor molecules are released, the camphor-rich area (the one without tracer particles) grows large and contacts with almost a half of the droplet periphery. At this stage, the droplet acceleration reaches the maximum (Figs. 11(c)–(d)). The droplet moves to the area with low camphor concentration and, as a consequence, the gradient of surface tension becomes low. In this phase of motion, the droplet slowed down and returned to the disk shape as shown in Figs. 11(e)–(f).
Figure 11: Snapshots of floating small particles during an intermittent motion of a paraffin droplet (volume $20~{}\mu\mathrm{L}$) containing camphor (concentration $0.06~{}\mathrm{M}$) and oil red O (concentration $0.8~{}\mathrm{g/L}$) illustrating the surface flows around the droplet via locations of tracer particles. The snapshot (a) was made $1500~{}\mathrm{s}$ after placing the droplet on the water surface. The time difference between subsequent snapshots is $2~{}\mathrm{s}$.
In the experiments, the continuous motion changed to the intermittent motion as time passed as shown in Fig. 1. The droplet motion should be initiated by a local fluctuations in camphor surface concentration. Once such asymmetric concentration drives the droplet, the droplet keeps moving in the initial stage. In contrast, in the latter stage, the droplet cannot keep moving but shows an intermittent motion. This should be because of the less camphor release rate due to the decrease in camphor concentration in the droplet. However, the paraffin droplet with less camphor concentration did not show intermittent motion. This implies the modification of water-oil interface by the presence of oil red O molecules. The release rate may be affected by the flow around the droplet[45]. For the clarification of the more detailed mechanism of this droplet deformation phenomenon, we need more precise measurement of the dynamics of camphor and oil red O molecules. We also need to construct a mathematical model that reproduces the time evolution of motion and deformation including hydrodynamics in future study.
@@ V Conclusions
We experimentally observed that a paraffin droplet containing camphor and oil red O exhibits spontaneous motion with deformation. At considered concentrations of camphor and oil red O, the droplet started to perform intermittent motion approximately $2000~{}\mathrm{s}$ after placing it on the water surface. We defined the variables characterizing large deformation and evaluated the relationships between the velocity and deformation magnitude. We also measured the time change in oil/water interfacial tension and found that it gradually decreased in time due to the adsorption of oil red O at the oil/water interface. Simultaneously, the camphor concentration in the droplet should also decrease with time, resulting in a smaller driving force. The combination of decreased interfacial tension due to oil red O and reduced driving force due to reduced camphor concentration in the droplet causes intermittent motion with dynamic deformation into a crescent shape.
@@ A.2 Analysis method
A.3 Derivation of the expression for line tension
Assuming the characteristic length of deformation of a triple line is much larger than the capillary length, we consider the deformation of the interface near the straight triple line. We set the Cartesian coordinates so that the $y$-axis matches the triple line and $z=\zeta_{\mathrm{ao}}(x)\,(x>0)$ and $z=-\zeta_{\mathrm{ow}}(x)\,(x>0)$ indicate the interface between air and oil and that between oil and water, respectively, as shown in Fig. 8. Then, the excess energy of the triple line per unit length is described as
@@ A.3 Derivation of the expression for line tension
$\qquad+\frac{1}{2}\rho_{\mathrm{ow}}g(e_{\mathrm{ow}}-\zeta_{\mathrm{ow}}(x))^{2}\biggr{\}},$ (18)
@@
where $\kappa_{\mathrm{ao}}^{-1}$ and $\kappa_{\mathrm{ow}}^{-1}$ are capillary lengths of air/oil and oil/water interfaces, respectively. From these equations, we derived Eq. (3).
Appendix B Water surface tension measurement around a droplet
The surface tension around a $50~{}\mu\mathrm{L}$ paraffin droplet with camphor (concentration $0.06~{}\mathrm{M}$) and oil red O (concentration $1.0~{}\mathrm{g/L}$) was measured using quasi-elastic laser scattering (QELS) method[8]. We put a metal screw on the bottom of the Petri dish upside down and then placed a droplet on the tip of the screw so that the droplet was pinned. We measured the surface tension at the $5~{}\mathrm{mm}$ distance from the droplet. The optical settings for the surface tension measurement were the same as described in [6, 7].
The surface tension $\gamma_{\mathrm{aw}}$ as the function of time $t$ is plotted in Fig. 13. Time $t=0$ corresponds to the moment when the droplet was placed on the tip of the screw. The average values of the surface tension $\gamma_{\mathrm{aw}}$ were $72.70~{}\mathrm{mN/m}$ at $-4~{}\mathrm{s}<t<-1~{}\mathrm{s}$ and $72.45~{}\mathrm{mN/m}$ at $1~{}\mathrm{s}<t<4~{}\mathrm{s}$. Using these values, we estimate $\delta\gamma_{\mathrm{aw}}$ as $0.25~{}\mathrm{mN/m}$ for the discussion in subsection IV.1.
Figure 13: (a) Geometry of experimental setup. (b) Surface tension as a function of time $t$ after locating a droplet on the water surface. The red shaded regions mark time intervals $-4~{}\mathrm{s}<t<-1~{}\mathrm{s}$ and $1~{}\mathrm{s}<t<4~{}\mathrm{s}$. The paraffin droplet included camphor at $0.06~{}\mathrm{M}$ and oil red O at $1.0~{}\mathrm{g/L}$.